the area of a triangle is 30 in. The base of the triangle measures 2in more than twice the height of the triangle. Find the measures of the base and height
Answer by Boreal(15235)

b=base
2h=b-2
h=(b-2)/2
A=(1/2)bh
=b*(b-2)/4=30
(b^2-2b)/4=30
b^2-2b=120
b^2-2b-120=0
(b-12)(b+10)=0
b=12 in
h=5 in
A=(1/2)bh=(1/2)*60=30 in^2